OpenStudy (anonymous):

Explain How You Can Use A Speedometer And A Clock To Tell How Far You Have Traveled In A Car If The Car's Odometer Is Not Working.(HINT: Assume You Are Traveling At A Constant Velocity.)please help and i will be ur fan?!?!?!:)

OpenStudy (danjs):

the speedometer will tell you Distance / time the clock will tell you time

OpenStudy (danjs):

\[\frac{ distance }{ time }*\frac{ time }{ 1 } = distance\]

OpenStudy (danjs):

so what will you do with the 2 numbers from the data you collected in order to get the distance traveled

OpenStudy (anonymous):

But what numbers?

OpenStudy (danjs):

it doesn't matter here, they just want to know how you would use the data , to figure out the distance

OpenStudy (danjs):

say you look at the speedometer and see X mi/h and you look at the clock and read traveled for Y hours \[\frac{ X ~miles }{ hour }*\frac{ Y~Hours }{ 1 } = (X*Y) ~~miles\]

OpenStudy (danjs):

Distance = Rate x time

OpenStudy (anonymous):

thank u so so so so so so much!!!!!!!!!!!!!!!!!!